# Ortho-Radial Drawing in Near-Linear Time

### v2

@@ 5 Extracting a strictly monotone cycle
We analyze the structural properties of the edges $(e_{1},e_{2},\ldots,e_{s})$ and the their incident faces $F_{i,i+1}$. The following lemma proves the intuitive fact that the rotation from the reference edge $e^{\star}$ to any $\overline{e_{i}}$ must be $90^{\circ}$ via any crossing-free path $P$ in $G_{k}^{+}$. Note that such a path $P$ must exist.
@@
Proof.
See the left drawing of Figure 12 for an illustration of the proof. Consider the ortho-radial representation $\mathcal{R}^{\prime}$ resulting from connecting the south endpoints $x_{1},x_{2},\ldots,x_{s}$ of the edges $e_{1},e_{2},\ldots,e_{s}$ in $G_{k}^{+}$ into a cycle $C=(x_{1},x_{2},\ldots,x_{s})$ in such a way that the rotation from $\overline{e_{i}}$ to the new edge $(x_{i},x_{i+1})$ is a $90^{\circ}$ left turn, the rotation from $\overline{e_{i}}$ to the new edge $(x_{i},x_{i-1})$ is a $90^{\circ}$ right turn, and any subpath of $C$ is a straight line. Observe that $C$ is a horizontal segment and is the facial cycle of the central face of $\mathcal{R}^{\prime}$. By (D1) and (D2), it is straightforward to convert a good drawing of $G_{k}$ into an ortho-radial drawing of $\mathcal{R}^{\prime}$, so $\mathcal{R}^{\prime}$ is drawable. Since $C$ is a horizontal segment, the edge label $\ell_{C}(e)$ of all edges $e\in C$ must be the same. Since $C$ cannot be a strictly monotone cycle, the only possibility is that $C$ is a monotone cycle, meaning that $\ell_{C}(e)=0$ for all edges $e\in C$. Now consider the path $P$ in the lemma statement and the edge $e=(x_{i},x_{i+1})$ in $C$. Since $\ell_{C}(e)=0$, we have $\mathsf{rotation}(P\circ e)=0$. Since $P\circ e$ makes a $90^{\circ}$ left turn at $x_{i}$, we have $\mathsf{rotation}(P)=1$. ∎
@@
Proof.
See the middle drawing of Figure 12 for an illustration of the proof. Consider the ortho-radial representation $\mathcal{R}^{\prime}$ resulting from connecting the south endpoints $x_{i}$ and $x_{j}$ of the edges $e_{i}$ and $e_{j}$ in $G_{k}^{+}$ into a horizontal edge $e^{\prime}=(x_{j},x_{i})$ in such a way that the path $\overline{e_{j}}\circ e^{\prime}\circ e_{i}$ makes two $90^{\circ}$ right turns. Similar to the proof of Lemma 5.1, $\mathcal{R}^{\prime}$ is drawable by extending a good drawing of $G_{k}$. The path $P^{\mathsf{out}}_{i\rightarrow j}$ together with the path $\overline{e_{j}}\circ e^{\prime}\circ e_{i}$ forms a facial cycle of a regular face. By (R2), we must have $\mathsf{rotation}(P^{\mathsf{out}}_{i\rightarrow j})=2$, as the sum of rotations for a regular face has to be $4$. ∎
Figure 12: Illustration for the proof of Lemmas 5.1, 5.2 and 5.3.
@@
Proof.
See the right drawing of Figure 12 for an illustration of the proof. Consider subgraph $H$ of $\tilde{G}$ induced by $G_{k}^{+}$ and all edges in $P$. Observe that $C$ is a facial cycle of $H$. If the central face of $\tilde{G}$ lies in the interior of $C$, then $C$ is the facial cycle of the central face of $H$, so $\mathsf{rotation}(C)=0$ by (R2). By Lemma 5.2, $\mathsf{rotation}(P^{\mathsf{out}}_{i\rightarrow j})=2$, so we must have $\mathsf{rotation}(P)=-2$. If the central face of $\tilde{G}$ lies in the exterior of $C$, then $C$ is the facial cycle of a regular face of $H$, so $\mathsf{rotation}(C)=4$ by (R2). Therefore, Lemma 5.2 implies that $\mathsf{rotation}(P)=2$. ∎
@@
Proof.
See the left drawing of Figure 13 for an illustration of the proof. Consider the path $P^{\mathsf{out}}_{i\rightarrow i+1}$, which is a subpath of $C$ starting at $e_{i}$ and ending at $\overline{e_{i+1}}$. By Lemma 5.2, $\mathsf{rotation}(P^{\mathsf{out}}_{i\rightarrow i+1})=2$, so there exists an edge $e^{\prime}$ in $P^{\mathsf{out}}_{i\rightarrow i+1}$ such that $\mathsf{rotation}(e_{i}\circ\cdots\circ e^{\prime})=1$ and $\mathsf{rotation}(e^{\prime}\circ\cdots\circ\overline{e_{i+1}})=1$ along the path $P^{\mathsf{out}}_{i\rightarrow i+1}$, or equivalently along the cycle $C$. Since $e_{i}$ and $e_{i+1}$ are vertical, such an edge $e^{\prime}$ must be horizontal. Since $e^{\prime}$ is in $G_{k}$, $e^{\prime}$ belongs to some horizontal segment $S_{l}$ in $A=(S_{1},S_{2},\ldots,S_{k})$. ∎
@@
The proof for the second statement of the lemma is similar. Suppose that $F_{i,i+1}$ is of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ) and some number $z_{l}$ in the string $Z_{i\rightarrow i+1}$ is at least 00 and $l\neq 1$. Then we can find a substring $(-2)\circ(-1)^{c}\circ 0$, for some $c\geq 1$, of $Z_{i\rightarrow i+1}$, and then we obtain a contradiction, as the horizontal segment corresponding to the substring $(-1)^{c}$ cannot exist. ∎
Figure 13: Illustration for the proof of Lemmas 5.4 and 5.5.
Intuitively, if a face $F_{i,i+1}$ is of type (,)square-unionsquare-union($\sqcup$,$\sqcup$)( ⊔ , ⊔ ), then a $\sqcap$-shape must exist in the middle of $P_{i\leftarrow i+1}$, and the horizontal segment corresponding to the middle part of the $\sqcap$-shape must be eligible for adding a virtual edge, so $F_{i,i+1}$ cannot be of type (,)square-unionsquare-union($\sqcup$,$\sqcup$)( ⊔ , ⊔ ). In the following lemma, we prove this intuitive observation formally, by combining Lemmas 5.3 and 5.5.
@@
To finish the proof, we just need to show that when $F_{i,i+1}$ is of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ), $F_{i,i+1}$ also has to be a regular face. Again, Lemma 5.3 implies that if $F_{i,i+1}$ is the central face, then the rotation of the path $P_{i\rightarrow i+1}=\overline{P_{i\leftarrow i+1}}$ is $2$. This contradicts Lemma 5.5, since it requires the rotation of $P_{i\rightarrow i+1}$ to be at most $-1$. Therefore, $F_{i,i+1}$ is a regular face. ∎
As discussed in the previous section, we may assume that each vertex in $\tilde{G}$ is incident to a horizontal segment. Consider the horizontal segment $S$ incident to the south endpoint $x_{i}$ of some $e_{i}\in\mathcal{N}_{\mathsf{south}}(A)$. In view of (S2), there are two possible reasons for why adding $S$ to the current good sequence $A$ does not result in a good sequence. The first possible reason is that $\mathcal{N}_{\mathsf{north}}(S)$ contains an edge that is not in $\mathcal{N}_{\mathsf{south}}(A)$. The second possible reason is that there exist two edges $e$ and $e^{\prime}$ such that $e^{\prime}$ immediately follows $e$ in the ordering of $\mathcal{N}_{\mathsf{north}}(S)$ and $e^{\prime}$ does not immediately follows $e$ in the ordering of $\mathcal{N}_{\mathsf{south}}(A)$. We show that the second reason is not possible.
@@
Our choice of $e_{i}$, $e_{j}$, and $S$ implies that for each horizontal segment $S^{\prime}$ such that $\mathcal{N}_{\mathsf{north}}(S^{\prime})$ contains an edge in $(e_{i+1},e_{i+2},\ldots,e_{j-1})$, the intersection of $\mathcal{N}_{\mathsf{north}}(S^{\prime})$ and $\mathcal{N}_{\mathsf{south}}(A)$ must be a contiguous subsequence of $(e_{i+1},e_{i+2},\ldots,e_{j-1})$, since otherwise we should select $S^{\prime}$ and not select $S$. We partition $(e_{i+1},e_{i+2},\ldots,e_{j-1})$ into groups according to the horizontal segment $S^{\prime}$ incident to the the south endpoint $x_{l}$ of each edge $e_{l}=(x_{l},y_{l})$. In other words, if the south endpoints of two edges of $(e_{i+1},e_{i+2},\ldots,e_{j-1})$ are both incident to the same horizontal segment $S^{\prime}$, then these two edges are in the same group corresponding to $S^{\prime}$. Note that each group corresponds to a contiguous subsequence of $(e_{i+1},e_{i+2},\ldots,e_{j-1})$.
@@
By Lemma 5.6, $F_{i,i+1}=F_{c_{1},c_{1}+1}$ cannot be also of type (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ), as this forces the face to be of type (,)square-unionsquare-union($\sqcup$,$\sqcup$)( ⊔ , ⊔ ). In view of the discussion above, the fact that $F_{c_{1},c_{1}+1}$ cannot be of type (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ) forces the type of $F_{c_{2},c_{2}+1}$ to be (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ). Similarly, we can argue that the types of $F_{c_{3},c_{3}+1},F_{c_{4},c_{4}+1},\ldots$ must be (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ), implying that the type of $F_{c_{t},c_{t}+1}=F_{j-1,j}$ is also (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ). The same argument for showing that $F_{i,i+1}$ is of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ) can also be used to show that $F_{j-1,j}$ must be of type (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ). Therefore, $F_{j-1,j}$ is of type (,)square-unionsquare-union($\sqcup$,$\sqcup$)( ⊔ , ⊔ ), which is impossible due to Lemma 5.6, so the lemma statement holds. ∎
Figure 14: Illustration for the proof of Lemmas 5.7 and 5.8.
@@
By Lemma 5.6, the face $F_{a_{i}-1,a_{i}}$ cannot be of type (,)square-unionsquare-union($\sqcup$,$\sqcup$)( ⊔ , ⊔ ), for each $1\leq i\leq t$. Therefore, the only possibility is that they are all of type (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ) or they are all of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ).
Now consider any face $F_{l,l+1}$ that is not of the form $F_{a_{i}-1,a_{i}}$ for some $1\leq i\leq t$. That is, $e_{l}$ and $e_{l+1}$ are in the same group, meaning that their south endpoints are both incident to the same horizontal segment $S^{\prime}$, and $e_{l+1}$ immediately follows $e_{l}$ in the sequential ordering $\mathcal{N}_{\mathsf{north}}(S^{\prime})$, so the face $F_{l,l+1}$ must be of type ()(-)( - ). ∎
@@
Proof.
If $F_{i,i+1}$ is of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ) or (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ), then Lemma 5.6 implies that $F_{i,i+1}$ is a regular face. By Lemma 5.7, the only remaining case is when $F_{i,i+1}$ is of type ()(-)( - ). Suppose that $F_{i,i+1}$ is of type ()(-)( - ), and let $C$ be the facial cycle of $F_{i,i+1}$. Then $\mathsf{rotation}(C)$ equals the sum of $\mathsf{rotation}(P^{\mathsf{out}}_{i\rightarrow i+1})$ and $\mathsf{rotation}(P_{i\leftarrow i+1})$. By the definition of the type ()(-)( - ), we know that $\mathsf{rotation}(P_{i\leftarrow i+1})=2$. By Lemma 5.2, we know that $\mathsf{rotation}(P^{\mathsf{out}}_{i\rightarrow i+1})=2$. Therefore, $\mathsf{rotation}(C)=4$, so $C$ is a regular face in view of (R2). ∎
@@
We define $P_{1}$ to be the subpath of $C$ starting at $\overline{e_{i^{\prime}+1}}$ and ending at $e_{i}$ and define $P_{2}$ to be the subpath of $C$ starting at $\overline{e_{i+1}}$ and ending at $e_{i^{\prime}}$. Therefore, $\mathsf{rotation}(C)$ is the sum of the rotation of the four paths $P^{\mathsf{out}}_{i\rightarrow i+1}$, $P^{\mathsf{out}}_{i^{\prime}\rightarrow i^{\prime}+1}$, $P_{1}$, and $P_{2}$. By Lemma 5.9, $F$ is a regular face, so $\mathsf{rotation}(C)=4$ due to (R2).
@@
By a similar argument that considers the suffix $P_{1}$ of the path $P_{i\leftarrow i+1}$, we may also force $F_{i,i+1}$ to be of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ). The fact that both types (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ) and (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ) exist is a contradiction to Lemma 5.8, so $F_{i,i+1}$ and $F_{i^{\prime},i^{\prime}+1}$ cannot be the same face. ∎
@@
Proof.
By Lemmas 5.9 and 5.10, the cycle formed by concatenating the paths $P_{i\rightarrow i+1}$, excluding the first edge $\overline{e_{i}}$ and the last edges $e_{i+1}$, over all $1\leq i\leq s$, separate the central face of $\tilde{G}$ from the outer face of $\tilde{G}$ and all faces $F_{i,i+1}$, for all $1\leq i\leq s$. Therefore, the central face and the outer face of $H$ are distinct, where the central face of $H$ contains the central face of $\tilde{G}$, and the outer face of $H$ contains the outer face of $\tilde{G}$ and also the faces $F_{i,i+1}$, for all $1\leq i\leq s$.
@@
Figure 16: Showing that $C$ is strictly monotone.
@@
To finish the proof, we still need to show that $C$ is strictly monotone. That is, we need to show that we cannot have $\ell_{C}(e)=0$ for all $e\in C$. Now suppose that $\ell_{C}(e)=0$ for all $e\in C$. We aim to show that this assumption implies that all faces $F_{i,i+1}$ are of the type ()(-)( - ), contradicting Lemma 5.8. See Figure 16 for an illustration of this part of the proof.
Again, here we assume that all faces $F_{i,i+1}$ are of the type ()(-)( - ) and (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ). The proof for the other case where all faces $F_{i,i+1}$ are of the type ()(-)( - ) and (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ) is similar. Consider any $e\in C$. We know that $e\in P_{i\rightarrow i+1}$ for some $1\leq i\leq s$. We extend $e$ into a subpath $P$ of $C$ such that $P$ is a subpath of $P_{i\rightarrow i+1}$, and both the edge immediately preceding $P$ and the edge immediately follows $P$ in $C$ are not in $P_{i\rightarrow i+1}$. Since the edge labels $\ell_{C}(\tilde{e})$ of all edges $\tilde{e}$ of $P$ are 00, the rotation from $\overline{e_{i}}$ to any edge $\tilde{e}$ of $P$ along the path $P_{i\rightarrow i+1}$ must be $-1$. By also considering the two edges of $P_{i\rightarrow i+1}$ immediately before and after $P$, we obtain a sequence of rotation numbers $0\circ(-1)^{c}\circ(-2)$ that is a substring of $Z_{i\rightarrow i+1}$. Such a substring corresponds to a $\sqcup$-shape in the drawing.
@@
We may repeat the same analysis for $F_{i-2,i-1}$, $F_{i-3,i-2}$, and so on, to infer that all of them are of type ()(-)( - ), contradicting Lemma 5.8, so we cannot have $\ell_{C}(e)=0$ for all $e\in C$. Therefore, $C$ is strictly monotone. ∎

### v3

@@ 5 Extracting a strictly monotone cycle
We analyze the structural properties of the edges $(e_{1},e_{2},\ldots,e_{s})$ and their incident faces $F_{i,i+1}$. The following lemma proves the intuitive fact that the rotation from the reference edge $e^{\star}$ to any $\overline{e_{i}}$ must be $90^{\circ}$ via any crossing-free path $P$ in $G_{k}^{+}$. Note that such a path $P$ must exist.
@@
Proof 5.2.
See the left drawing of Figure 12 for an illustration of the proof. Consider the ortho-radial representation $\mathcal{R}^{\prime}$ resulting from connecting the south endpoints $x_{1},x_{2},\ldots,x_{s}$ of the edges $e_{1},e_{2},\ldots,e_{s}$ in $G_{k}^{+}$ into a cycle $C=(x_{1},x_{2},\ldots,x_{s})$ in such a way that the rotation from $\overline{e_{i}}$ to the new edge $(x_{i},x_{i+1})$ is a $90^{\circ}$ left turn, the rotation from $\overline{e_{i}}$ to the new edge $(x_{i},x_{i-1})$ is a $90^{\circ}$ right turn, and any subpath of $C$ is a straight line. Observe that $C$ is a horizontal segment and is the facial cycle of the central face of $\mathcal{R}^{\prime}$. By 1 and 2, it is straightforward to convert a good drawing of $G_{k}$ into an ortho-radial drawing of $\mathcal{R}^{\prime}$, so $\mathcal{R}^{\prime}$ is drawable. Since $C$ is a horizontal segment, the edge label $\ell_{C}(e)$ of all edges $e\in C$ must be the same. Since $C$ cannot be a strictly monotone cycle, the only possibility is that $C$ is a monotone cycle, meaning that $\ell_{C}(e)=0$ for all edges $e\in C$. Now consider the path $P$ in the lemma statement and the edge $e=(x_{i},x_{i+1})$ in $C$. Since $\ell_{C}(e)=0$, we have $\mathsf{rotation}(P\circ e)=0$. Since $P\circ e$ makes a $90^{\circ}$ left turn at $x_{i}$, we have $\mathsf{rotation}(P)=1$.
@@
Proof 5.4.
See the middle drawing of Figure 12 for an illustration of the proof. Consider the ortho-radial representation $\mathcal{R}^{\prime}$ resulting from connecting the south endpoints $x_{i}$ and $x_{j}$ of the edges $e_{i}$ and $e_{j}$ in $G_{k}^{+}$ into a horizontal edge $e^{\prime}=(x_{j},x_{i})$ in such a way that the path $\overline{e_{j}}\circ e^{\prime}\circ e_{i}$ makes two $90^{\circ}$ right turns. Similar to the proof of Lemma 5.1, $\mathcal{R}^{\prime}$ is drawable by extending a good drawing of $G_{k}$. The path $P^{\mathsf{out}}_{i\rightarrow j}$ together with the path $\overline{e_{j}}\circ e^{\prime}\circ e_{i}$ forms a facial cycle of a regular face. By 2, we must have $\mathsf{rotation}(P^{\mathsf{out}}_{i\rightarrow j})=2$, as the sum of rotations for a regular face has to be $4$.
Figure 12: Illustration for the proof of Lemmas 5.1, 5.3 and 5.5.
@@
Proof 5.6.
See the right drawing of Figure 12 for an illustration of the proof. Consider the subgraph $H$ of $\tilde{G}$ induced by $G_{k}^{+}$ and all edges in $P$. Observe that $C$ is a facial cycle of $H$. If the central face of $\tilde{G}$ lies in the interior of $C$, then $C$ is the facial cycle of the central face of $H$, so $\mathsf{rotation}(C)=0$ by 2. By Lemma 5.3, $\mathsf{rotation}(P^{\mathsf{out}}_{i\rightarrow j})=2$, so we must have $\mathsf{rotation}(P)=-2$. If the central face of $\tilde{G}$ lies in the exterior of $C$, then $C$ is the facial cycle of a regular face of $H$, so $\mathsf{rotation}(C)=4$ by 2. Therefore, Lemma 5.3 implies that $\mathsf{rotation}(P)=2$.
@@
Proof 5.8.
See the left drawing of Figure 13 for an illustration of the proof. Consider the path $P^{\mathsf{out}}_{i\rightarrow i+1}$, which is a subpath of $C$ starting at $e_{i}$ and ending at $\overline{e_{i+1}}$. By Lemma 5.3, $\mathsf{rotation}(P^{\mathsf{out}}_{i\rightarrow i+1})=2$, so there exists an edge $e^{\prime}$ in $P^{\mathsf{out}}_{i\rightarrow i+1}$ such that $\mathsf{rotation}(e_{i}\circ\cdots\circ e^{\prime})=1$ and $\mathsf{rotation}(e^{\prime}\circ\cdots\circ\overline{e_{i+1}})=1$ along the path $P^{\mathsf{out}}_{i\rightarrow i+1}$, or equivalently along the cycle $C$. Since $e_{i}$ and $e_{i+1}$ are vertical, such an edge $e^{\prime}$ must be horizontal. Since $e^{\prime}$ is in $G_{k}$, $e^{\prime}$ belongs to some horizontal segment $S_{l}$ in $A=(S_{1},S_{2},\ldots,S_{k})$.
@@
The proof for the second statement of the lemma is similar. Suppose that $F_{i,i+1}$ is of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ) and some number $z_{l}$ in the string $Z_{i\rightarrow i+1}$ is at least 00 and $l\neq 1$. Then we can find a substring $(-2)\circ(-1)^{c}\circ 0$, for some $c\geq 1$, of $Z_{i\rightarrow i+1}$, and then we obtain a contradiction, as the horizontal segment corresponding to the substring $(-1)^{c}$ cannot exist.
Figure 13: Illustration for the proof of Lemmas 5.7 and 5.9.
Intuitively, if a face $F_{i,i+1}$ is of type (,)square-unionsquare-union($\sqcup$,$\sqcup$)( ⊔ , ⊔ ), then a $\sqcap$-shape must exist in the middle of $P_{i\leftarrow i+1}$, and the horizontal segment corresponding to the middle part of the $\sqcap$-shape must be eligible for adding a virtual edge, so $F_{i,i+1}$ cannot be of type (,)square-unionsquare-union($\sqcup$,$\sqcup$)( ⊔ , ⊔ ). In the following lemma, we prove this intuitive observation formally, by combining Lemmas 5.5 and 5.9.
@@
To finish the proof, we just need to show that when $F_{i,i+1}$ is of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ), $F_{i,i+1}$ also has to be a regular face. Again, Lemma 5.5 implies that if $F_{i,i+1}$ is the central face, then the rotation of the path $P_{i\rightarrow i+1}=\overline{P_{i\leftarrow i+1}}$ is $2$. This contradicts Lemma 5.9, since it requires the rotation of $P_{i\rightarrow i+1}$ to be at most $-1$. Therefore, $F_{i,i+1}$ is a regular face.
As discussed in the previous section, we may assume that each vertex in $\tilde{G}$ is incident to a horizontal segment. Consider the horizontal segment $S$ incident to the south endpoint $x_{i}$ of some $e_{i}\in\mathcal{N}_{\mathsf{south}}(A)$. In view of 2, there are two possible reasons for why adding $S$ to the current good sequence $A$ does not result in a good sequence. The first possible reason is that $\mathcal{N}_{\mathsf{north}}(S)$ contains an edge that is not in $\mathcal{N}_{\mathsf{south}}(A)$. The second possible reason is that there exist two edges $e$ and $e^{\prime}$ such that $e^{\prime}$ immediately follows $e$ in the ordering of $\mathcal{N}_{\mathsf{north}}(S)$ and $e^{\prime}$ does not immediately follow $e$ in the ordering of $\mathcal{N}_{\mathsf{south}}(A)$. We show that the second reason is not possible.
@@
Our choice of $e_{i}$, $e_{j}$, and $S$ implies that for each horizontal segment $S^{\prime}$ such that $\mathcal{N}_{\mathsf{north}}(S^{\prime})$ contains an edge in $(e_{i+1},e_{i+2},\ldots,e_{j-1})$, the intersection of $\mathcal{N}_{\mathsf{north}}(S^{\prime})$ and $\mathcal{N}_{\mathsf{south}}(A)$ must be a contiguous subsequence of $(e_{i+1},e_{i+2},\ldots,e_{j-1})$, since otherwise we should select $S^{\prime}$ and not select $S$. We partition $(e_{i+1},e_{i+2},\ldots,e_{j-1})$ into groups according to the horizontal segment $S^{\prime}$ incident to the south endpoint $x_{l}$ of each edge $e_{l}=(x_{l},y_{l})$. In other words, if the south endpoints of two edges of $(e_{i+1},e_{i+2},\ldots,e_{j-1})$ are both incident to the same horizontal segment $S^{\prime}$, then these two edges are in the same group corresponding to $S^{\prime}$. Note that each group corresponds to a contiguous subsequence of $(e_{i+1},e_{i+2},\ldots,e_{j-1})$.
@@
By Lemma 5.11, $F_{i,i+1}=F_{c_{1},c_{1}+1}$ cannot be also of type (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ), as this forces the face to be of type (,)square-unionsquare-union($\sqcup$,$\sqcup$)( ⊔ , ⊔ ). In view of the discussion above, the fact that $F_{c_{1},c_{1}+1}$ cannot be of type (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ) forces the type of $F_{c_{2},c_{2}+1}$ to be (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ). Similarly, we can argue that the types of $F_{c_{3},c_{3}+1},F_{c_{4},c_{4}+1},\ldots$ must be (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ), implying that the type of $F_{c_{t},c_{t}+1}=F_{j-1,j}$ is also (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ). The same argument for showing that $F_{i,i+1}$ is of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ) can also be used to show that $F_{j-1,j}$ must be of type (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ). Therefore, $F_{j-1,j}$ is of type (,)square-unionsquare-union($\sqcup$,$\sqcup$)( ⊔ , ⊔ ), which is impossible due to Lemma 5.11, so the lemma statement holds.
Figure 14: Illustration for the proof of Lemmas 5.13 and 5.15.
@@
By Lemma 5.11, for each $1\leq i\leq t$, the face $F_{a_{i}-1,a_{i}}$ cannot be of type (,)square-unionsquare-union($\sqcup$,$\sqcup$)( ⊔ , ⊔ ). Therefore, the only possibility is that they are all of type (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ) or all of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ).
Now consider any face $F_{l,l+1}$ that is not of the form $F_{a_{i}-1,a_{i}}$ for some $1\leq i\leq t$. That is, $e_{l}$ and $e_{l+1}$ are in the same group, meaning that their south endpoints are both incident to the same horizontal segment $S^{\prime}$, and $e_{l+1}$ immediately follows $e_{l}$ in the sequential ordering $\mathcal{N}_{\mathsf{north}}(S^{\prime})$, so the face $F_{l,l+1}$ must be of type ()(-)( - ).
@@
Proof 5.18.
If $F_{i,i+1}$ is of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ) or (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ), then Lemma 5.11 implies that $F_{i,i+1}$ is a regular face. By Lemma 5.13, the only remaining case is when $F_{i,i+1}$ is of type ()(-)( - ). Suppose that $F_{i,i+1}$ is of type ()(-)( - ), and let $C$ be the facial cycle of $F_{i,i+1}$. Then $\mathsf{rotation}(C)$ equals the sum of $\mathsf{rotation}(P^{\mathsf{out}}_{i\rightarrow i+1})$ and $\mathsf{rotation}(P_{i\leftarrow i+1})$. By the definition of the type ()(-)( - ), we know that $\mathsf{rotation}(P_{i\leftarrow i+1})=2$. By Lemma 5.3, we know that $\mathsf{rotation}(P^{\mathsf{out}}_{i\rightarrow i+1})=2$. Therefore, $\mathsf{rotation}(C)=4$, so $C$ is a regular face in view of 2.
@@
We define $P_{1}$ to be the subpath of $C$ starting at $\overline{e_{i^{\prime}+1}}$ and ending at $e_{i}$ and define $P_{2}$ to be the subpath of $C$ starting at $\overline{e_{i+1}}$ and ending at $e_{i^{\prime}}$. Therefore, $\mathsf{rotation}(C)$ is the sum of the rotation of the four paths $P^{\mathsf{out}}_{i\rightarrow i+1}$, $P^{\mathsf{out}}_{i^{\prime}\rightarrow i^{\prime}+1}$, $P_{1}$, and $P_{2}$. By Lemma 5.17, $F$ is a regular face, so $\mathsf{rotation}(C)=4$ due to 2.
@@
By a similar argument that considers the suffix $P_{1}$ of the path $P_{i\leftarrow i+1}$, we may also force $F_{i,i+1}$ to be of type (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ). The fact that both types (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ) and (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ) exist is a contradiction to Lemma 5.15, so $F_{i,i+1}$ and $F_{i^{\prime},i^{\prime}+1}$ cannot be the same face.
@@
Proof 5.22.
By Lemmas 5.17 and 5.19, the cycle formed by concatenating the paths $P_{i\rightarrow i+1}$, excluding the first edge $\overline{e_{i}}$ and the last edges $e_{i+1}$, over all $1\leq i\leq s$, separate the central face of $\tilde{G}$ from the outer face of $\tilde{G}$ and all faces $F_{i,i+1}$, for all $1\leq i\leq s$. Therefore, the central face and the outer face of $H$ are distinct, where the central face of $H$ contains the central face of $\tilde{G}$, and the outer face of $H$ contains the outer face of $\tilde{G}$ and also the faces $F_{i,i+1}$, for all $1\leq i\leq s$.
@@
{sidefigure}
[t!] [Uncaptioned image] Showing that $C$ is strictly monotone.
@@
To finish the proof, we still need to show that $C$ is strictly monotone. That is, we need to show that we cannot have $\ell_{C}(e)=0$ for all $e\in C$. Now suppose that $\ell_{C}(e)=0$ for all $e\in C$. We aim to show that this assumption implies that all faces $F_{i,i+1}$ are of the type ()(-)( - ), contradicting Lemma 5.15. See Section 5 for an illustration of this part of the proof.
Again, here we assume that all faces $F_{i,i+1}$ are of the type ()(-)( - ) and (,)square-union($\sqcup$,$\ast$)( ⊔ , ∗ ). The proof for the other case where all faces $F_{i,i+1}$ are of the type ()(-)( - ) and (,)square-union($\ast$,$\sqcup$)( ∗ , ⊔ ) is similar. Consider any $e\in C$. We know that $e\in P_{i\rightarrow i+1}$ for some $1\leq i\leq s$. We extend $e$ into a subpath $P$ of $C$ such that $P$ is a subpath of $P_{i\rightarrow i+1}$, and both the edge immediately preceding $P$ and the edge immediately following $P$ in $C$ are not in $P_{i\rightarrow i+1}$. Since the edge labels $\ell_{C}(\tilde{e})$ of all edges $\tilde{e}$ of $P$ are 00, the rotation from $\overline{e_{i}}$ to any edge $\tilde{e}$ of $P$ along the path $P_{i\rightarrow i+1}$ must be $-1$. By also considering the two edges of $P_{i\rightarrow i+1}$ immediately before and after $P$, we obtain a sequence of rotation numbers $0\circ(-1)^{c}\circ(-2)$ that is a substring of $Z_{i\rightarrow i+1}$. Such a substring corresponds to a $\sqcup$-shape in the drawing.
@@
We may repeat the same analysis for $F_{i-2,i-1}$, $F_{i-3,i-2}$, and so on, to infer that all of them are of type ()(-)( - ), contradicting Lemma 5.15, so we cannot have $\ell_{C}(e)=0$ for all $e\in C$. Therefore, $C$ is strictly monotone.